# FLASC: A Flare-Sensitive Clustering Algorithm

### v1

@@ 4.3.2. Evaluation and settings
Finally, we selected the parameter values that maximised the average ARI and minimised the average centroid spread over all data sets. These parameter values were: min branch size $m_{b}=12$, the full cluster approximation graph, and eom branch selection strategy.
Figure 7. An illustration explaining the centroid spread stability metric. (a) A single point cloud with branch length $b_{l}=65$ and noise ratio $n_{r}=0.852$ coloured by the ground truth labels. The average position per ground truth label—i.e., true centroid—is indicated by black crosses: one for each branch and one for the entire cluster. (b) The weighted average position of each subgroup detected by FLASC—i.e., found centroids—for all $10$ data sets with these parameters coloured by the closest true centroid. (c) The unweighted average found centroid positions per subgroup shown as black pentagons. The centroid spread is quantified by the distances from the centroids to this root position normalised by the branch length $b_{l}$. Consequently, the metric describes stability, not accuracy.
A visual explanation of the centroid spread stability metric.
Figure 7. An illustration explaining the centroid spread stability metric. (a) A single point cloud with branch length $b_{l}=65$ and noise ratio $n_{r}=0.852$ coloured by the ground truth labels. The average position per ground truth label—i.e., true centroid—is indicated by black crosses: one for each branch and one for the entire cluster. (b) The weighted average position of each subgroup detected by FLASC—i.e., found centroids—for all $10$ data sets with these parameters coloured by the closest true centroid. (c) The unweighted average found centroid positions per subgroup shown as black pentagons. The centroid spread is quantified by the distances from the centroids to this root position normalised by the branch length $b_{l}$. Consequently, the metric describes stability, not accuracy.
4.3.3. Results
Figure 8. Data sets with varying noise ratios and branch lengths coloured by the FLASC labels. Notice that the central subgroup in red increases with the noise ratio. Branch lengths (columns) are normalised visually.
Scatterplots for the tree-pointed star point clouds for all branch length and noise ratio combinations with FLASC’s labelling indicating that FLASC detects the branches as a single subgroup in almost all combinations.
Figure 8. Data sets with varying noise ratios and branch lengths coloured by the FLASC labels. Notice that the central subgroup in red increases with the noise ratio. Branch lengths (columns) are normalised visually.
Figure 9. Heatmaps showing FLASC’s (a) average Adjusted Rand Index (ARI) and (b) $95$th percentile centroid spread over $10$ point-clouds generated with noise ratio $n_{r}$ and branch length $b_{l}$. The centroid spread distances were normalised with the branch length $b_{l}$, allowing for comparison between data sets with different parameters.
Two heatmaps indicating for which branch lengths and noise ratios FLASC performs well.
Figure 9. Heatmaps showing FLASC’s (a) average Adjusted Rand Index (ARI) and (b) $95$th percentile centroid spread over $10$ point-clouds generated with noise ratio $n_{r}$ and branch length $b_{l}$. The centroid spread distances were normalised with the branch length $b_{l}$, allowing for comparison between data sets with different parameters.
Figure 8 shows one data set for each branch length $b_{l}$ and noise ratio $n_{r}$ combination, where the data points are coloured by their FLASC labels. As expected, FLASC detected the three branches and the cluster’s centre in almost all data sets. This pattern was consistent across all $10$ data sets generated for each branch length $b_{l}$ and noise ratio $n_{r}$ combination, as shown by the average ARI value heatmap in Figure 9. This latter figure indicates that FLASC’s average ARI decreases as the noise ratio increases, which can be explained by an increasing number of points assigned to the centre label. While this centre label makes sense semantically, it is not present in the ground truth labels. Therefore, the ARI value is expected to decrease as more points are given the centre label. For this reason, it could be argued that the ARI values give a too-pessimistic view of FLASC’s performance on this data set.
FLASC’s stability was analysed through its centroid spread over different data set samples with the same branch length $b_{l}$ and noise ratio $n_{r}$. Figure 9 visualises the $95$th percentile of the centroid spread distances as a heatmap. In general, the centroid spread was relatively low, with the $95$th distance percentile being less than $\frac{1}{8}b_{l}$ in most of the evaluated noise ratio–branch length space. All in all, these values demonstrate that FLASC produces a similar output on data set samples with the same underlying distribution.
4.4. Computational Performance
@@
In this section, we explore how the computational cost of FLASC compares to other clustering algorithms. Given the challenges in accurately benchmarking the computational performance of algorithms (Kriegel et al., 2017), we limit this comparison to the trends in run time scaling over data set size and number of dimensions for specific implementations.
4.4.1. Datasets
@@
A Gaussian random walk process was used to generate data sets for the performance comparison. Specifically, for a space with $d$ dimensions, $c$ uniform random starting points were sampled in a volume that fits five times the number of to-be-generated clusters. Then, $5$ random walks, with $50$ steps each, were sampled from each starting point. Every step moved along one of the dimensions with a length sampled from a normal distribution ($\mu=0$, $\sigma=0.1$). The resulting point clouds have more varied properties than the Gaussian blobs McInnes and Healy (McInnes and Healy, 2017) used for their run time comparison of HDBSCAN*, including non-trivially varying densities and branching patterns within clusters. Note that the number of (density-based) clusters in each point cloud may differ from the number of starting points $c$ due to possible overlaps or sparse regions in the random walks.
4.4.2. Evaluation and settings
@@
The random walk data sets were generated with varying numbers of dimensions ($2$, $8$, $16$) and starting points ($2$ to $800$ in $10$ exponentially spaced steps). $10$ data sets were sampled for each combination of parameters, resulting in a total of $300$ point clouds.
@@
Three clustering algorithms’ Python implementations were compared: fastcluster (Müllner, 2013) (version 1.1.26), McInnes and Healy’s HDBSCAN* (McInnes et al., 2017) (version 0.8.28), and FLASC with the full and core approximation graphs. HDBSCAN* and both FLASC versions were evaluated with min samples $k=10$, min cluster size $m_{c}=100$, min branch size $m_{b}=20$, allow single cluster enabled, and their multiprocessing support disabled to better describe the algorithms’ intrinsic complexity. Fastcluster’s default parameter values were used, resulting in a single linkage dendrogram.
@@
Time measurements were performed on a computer with a $3.4$GHz Intel Core i7-6700 processor and $16$GB RAM. Each algorithm was evaluated on each data set once, recording the run time and number of detected clusters. The smallest data set for which an algorithm required more than $80$ seconds was recorded for each number of dimensions. Larger data sets were not evaluated for those algorithms.
Figure 10. Average run time in seconds over $10$ random walk point clouds with (a) 2 dimensions, (b) 8 dimensions, and (c) 16 dimensions. The shaded areas around each line indicate the mean’s $95\%$ confidence interval. Note the changes in the $y$ range between the sub-figures.
@@
Three line plots indicating the average run times of fastcluster, HDBSCAN*, and the two FLASC variants for random walk data sets with 2, 8, and 16 dimensions.
Figure 10. Average run time in seconds over $10$ random walk point clouds with (a) 2 dimensions, (b) 8 dimensions, and (c) 16 dimensions. The shaded areas around each line indicate the mean’s $95\%$ confidence interval. Note the changes in the $y$ range between the sub-figures.
4.4.3. Results
@@
Figure 10 shows the algorithms’ average run times in seconds over the data set size and number of dimensions. There are three patterns of note: firstly, fastcluster shows a quadratic trend, resulting in the longest run times. The other algorithms show sub-quadratic trends on the $2$ and $8$ dimensional data sets but approach fastcluster’s quadratic trend in the $16$ dimensional case. This is expected because constructing and querying space trees becomes more expensive with more dimensions. Secondly, HDBSCAN* and both FLASC variants scale similarly in all three conditions. HDBSCAN* is consistently the fastest of the three, followed by FLASC with the core approximation graph. However, their run time differences diminish in the higher dimensional cases, indicating that the additional cost of branch detection is relatively low compared to the cost of detecting the clusters.
@@
In conclusion, both FLASC variants’ computational performance scale similarly to HDBSCAN* and the more dimensions the data contains, the smaller the scaling trend differences between the algorithm.
5. Discussion
Two case studies on real-world data and two analyses on synthetic data were performed to demonstrate FLASC and its properties. Section 4.1 showed how the flare-detection post-processing step enabled the detection of subgroups representing two diabetes types. Section 4.2 showed a more complex data set in which the subgroups are detectable by both FLASC and HDBSCAN*. FLASC still provides a benefit for exploration because the uncovered branch hierarchy includes information about the connectivity between the subgroups. However, structure learning algorithms—such as Reversed Graph Embeddings (Mao et al., 2017) and Bonnaire et al.’s mixture model approach (Bonnaire et al., 2022)—can provide even more information about the data’s shape at more computational costs. Section 3.2 demonstrated FLASC’s stability by showing that its output is similar on multiple samples of the same underlying distribution. This analysis could be expanded to investigate how well FLASC deals with shapes with unequal branch lengths. The weighted average data point—and centrality metric as a result—may not accurately reflect the centre of such clusters. Consequently, FLASC’s branch hierarchy will be a less accurate representation of the underlying topology but should still detect the flares. Monocle 3 (Qiu et al., 2017) deals with this problem by selecting the centre point in a projection manually (Pliner et al., 2022). Section 4.4 demonstrated that FLASC’s computational performance scales similar to HDBSCAN*. The scaling trends appeared to become more similar as the data contained more dimensions. However, neither algorithm was evaluated with multiprocessing enabled, which can introduce run-time differences in practical applications. In addition, extracting the full cluster approximation graph can be more expensive than reported, depending on the data’s characteristics.
5.1. Branch-detection’s practical value
@@
As briefly mentioned in Section 1.2, the argument that branches are not detectable as clusters only applies when they do not contain local density maxima. In other cases, the subgroups can be detected as density-based clusters. If one is only interested in the existence of subgroups, then there may not be many data sets where FLASC provides a benefit over HDBSCAN*, as relevant (sparse) subgroups may be rare, and HDBSCAN*’s assumption that such points are noise may be valid. However, FLASC provides a benefit when looking for patterns in the relatively uncommon parts of a data set or when the cluster’s shapes are relevant. The algorithm, therefore, is not the default go-to solution for all things clustering. Instead, we envision it as a valuable tool for exploring unfamiliar data, providing guidance into which subpopulations exist and informing follow-up questions. Knowing that a cluster may represent multiple subgroups can be very relevant.
5.2. Alternative centrality metrics
@@
The presented FLASC algorithm uses a geometric distance-to-centroid centrality metric to describe how central data points lie within a cluster (Figure 3, Equation 3). An interesting alternative is a geodesic centrality, which measures the number of edges between each data point and the cluster’s root point in the cluster approximation graph. Here, the root point can be chosen as the data point closest to the cluster’s centroid, as we did for the branch-membership vectors in Section 3.1 (Figure 4). This geodesic centrality would agree with the notion that distances in high dimensional data may not accurately reflect distances along the intrinsic structure of a data set, which was one of the motivations for Reversed Graph Embeddings (Mao et al., 2017). It would also be closer to the maximum shortest-path centrality metric used by Vandaele et al. (Vandaele et al., 2021).
@@
Several trade-offs between the geometric and geodesic centrality metrics made us choose the geometric one. (1) Computing the geodesic centrality is more expensive because it requires an additional traversal over the entire cluster approximation graph. The extra cost, however, should be low compared to other parts of the algorithm. (2) The resolution of the geodesic centrality is lower, as it expresses the number of edges to the root point. As a result, zero-persistent branches are more likely to occur. In addition, it reduces the detectability of small branches that are well connected. On the other hand, that can be seen as beneficial noise suppression. In addition, the branch selection persistence parameter becomes more interpretable and would represent the traversal depth of a branch in the approximation graph. (3) The cluster’s centroid may lie outside of the cluster itself, resulting in a root point and centrality values that do not accurately describe its centre. For example, imagine a U-shaped cluster. The centroid would lie in between the two arms, and the root would lie in one of the arms. As a result, the geodesic metric would find one smaller and one larger branch rather than two equal branches. On the other hand, the geometric centrality finds the two branches and the connecting bend as three separate groups. Confusingly, it also contains two regions with a local centrality maximum, which FLASC gives a single label.
@@
FLASC’s general process can also be used with metrics that capture other aspects than data points’ centrality. At its core, FLASC consists of two filtrations, one to determine the connectivity between data points and one to analyse a signal on the resulting graph. The process would then describe how many distinct local minima (or maxima) of the metric exist within the clusters. The resulting interpretation does not have to relate to the cluster’s shape.
@@
One could even interpret FLASC as two applications of HDBSCAN*: one over the density and one over the eccentricity. This perspective raises a possible improvement to the algorithm by translating the mutual reachability concept to the centrality metric. The idea of ‘pushing away points in low-density regions’ can also be applied to the centrality and would emphasise the centrality difference between the centre and branch ends. Additionally, smoothing the centrality profile by incorporating neighbouring values could improve the algorithm’s robustness to noise. The additional computational cost should be low, as points’ neighbours are already known when the centrality is computed. Another way to improve noise robustness could be to implement the mutual $k$-nearest neighbour approach used by Dalmia and Sia (Dalmia and Sia, 2021) to improve UMAP projections. It would provide a subgraph of the core approximation graph that better reflects the cluster’s connectivity in high dimensional data sets. We leave evaluating these ideas for future work.
5.3. Visually summarising data’s shape
A strength of Mapper (Singh et al., 2007) and Reversed Graph Embeddings (Mao et al., 2017) is that they can summarise the data’s shape using intuitive visualisations. While FLASC’s branch-condensed tree provides some information about the clusters’ shapes, interpreting the cluster’s shape from it is not trivial. Practically, this means that detecting branch-based subgroups is simple, but interpreting trajectories or shapes is difficult. For example, in Section 4.2, some connections in the branch hierarchy (Figure 6(c)) are not apparent from the 2D densMAP (Narayan et al., 2021) projection (Figure 6(a)). However, such 2D projections should not be interpreted as a canonical representation of the data’s shape, as they do not necessarily retain the data’s local and global structures (Chari et al., 2021).
@@ 5.3. Visually summarising data’s shape
Studying how well two-dimensional layouts of FLASC’s cluster approximation graphs work as shape summarising visualisations would be an interesting future research direction. These graphs directly encode the connectivity used by the algorithm. Therefore, explicitly drawing their edges avoids the visual conflicts we found with the densMAP projections. Another benefit is that all (non-noise) data points are represented in the graphs once. Directly visualising the graphs, however, probably does not scale to larger sizes in terms of computational cost for the layout algorithm and visual interpretability. Ways to summarise the networks would have to be found, which could be based on $k$-means centroids like in Reversed Graph Embeddings or the local density maxima in the cluster.
6. Conclusion

### v2

@@ Introduction
Finally, we selected the parameter values that maximised the average ARI and minimised the average centroid spread over all data sets. For FLASC, these parameter values were: min branch size $m_{b}=12$, the core cluster approximation graph, and leaf branch selection strategy. Larger values for $m_{b}$ provided similar performance values, indicating the algorithm is not very sensitive to $m_{b}$ as long as its value is large enough to exclude small noisy structures. HDBSCAN* performed best with min cluster size $m_{c}=8$ and the leaf cluster selection strategy.
Results
(a) True and predicted subgroups
(b) Average ARI
(c) 95% centroid spread
Figure 5: Results for the stability benchmark. (a) One point cloud with $b_{l}=18$ and $n_{r}=0.47$ coloured by ground truth and predicted labels that characterise the algorithms’ behaviours. (b) Heatmap with average ARI values for FLASC and kMeans over all branch lengths and noise ratios. (c) Heatmap with 95% centroid spread values for FLASC and kMeans over all branch lengths and noise ratios.
Figure 5(a) demonstrates which subgroups the algorithms detect on a point cloud with $b_{l}=18$ and $n_{r}=0.47$. The true labels indicate that most confusion is expected in the dense centres, as the true labels are not spatially separable in those locations. Furthermore, FLASC neatly separated central points by the closest branch. kMeans, on the other hand, generally assigned all central points to one branch and limits the other branches to non-central points. Finally, as expected, HDBSCAN* did not detect the branch-based subgroups.
The patterns identified in the previous paragraph generalised to other branch lengths and noise ratios, as shown by the average ARI in Fig. 5(b). FLASC achieved ARI values $\geq 0.8$ on all branch lengths for low noise ratios. Its performance degrades as more noise is introduced, where the decrease in ARI is stronger for large branch lengths. This pattern can be explained by the confusion on central points, as there are more central points with longer branch lengths and larger noise ratios. kMeans achieved ARI values $\leq 0.8$. Its performance decreased most strongly with the branch length, which can be explained by the assignment of all central points to one branch.
The 95 percentile centroid spread, shown in Fig. 5(c), indicates FLASC provided more stable results than kMeans. Both algorithms behaved in a similar pattern: the spread was lowest with low noise values and small branch lengths. When either parameter increased, the spread increased as well. However, FLASC achieved lower spread values in these cases, mostly remaining below $0.25b_{l}$. For kMeans, the spread was larger than $0.25b_{l}$ in the upper $b_{l}$$n_{r}$ triangle. The behaviour on central points explains these patterns.
All in all, this benchmark showed that both FLASC and kMeans are able to detect the branch based subgroups. However, FLASC more accurately segmented the central points, which resulted in a higher accuracy and stability. In addition, FLASC achieved this result without being given the correct number of subgroups a priori.
Computational Performance
Now we turn to the computational performance. This second synthetic benchmark is designed to answer the following research question:
How does FLASC’s compute cost compare to HDBSCAN*, kMeans, and fastcluster?
kMeans is included in this comparison because it is a fast, popular clustering algorithm. HDBSCAN* is included to determine how much extra cost FLASC introduces. Fastcluster is included as an example of a quadratically scaling implementation (as shown in McInnes and Healy (2017)).
@@ Results
Given the challenges in accurately benchmarking the computational performance of algorithms (Kriegel et al., 2017), we limit this comparison to the trends in run time scaling over data set size and number of dimensions for specific implementations.
Datasets
@@
A Gaussian random walk process was used to generate data sets that contain non-trivially varying densities and branching structures in a controlled environment. For a space with $d$ dimensions, $c$ uniform random starting points were sampled in a volume that fits five times the number of to-be-generated clusters. Then, five $50$-step random walks were sampled from each starting point. Every step moved along one of the dimensions with a length sampled from a normal distribution ($\mu=0$, $\sigma=0.1$). The resulting point clouds have more varied properties than the Gaussian blobs used in McInnes and Healy (2017)’s run time comparison of HDBSCAN*. Note that the number of (density-based) clusters in each point cloud may differ from the number of starting points $c$ due to possible overlaps or sparse regions in the random walks.
Evaluation and settings
@@
The random walk data sets were generated with varying numbers of dimensions ($2$, $8$, $16$) and starting points ($2$ to $800$ in $10$ exponentially spaced steps). Ten data sets were sampled for each combination of parameters, resulting in a total of $300$ point clouds.
@@
The algorithms were compared using their Python implementations: fastcluster version 1.1.26 (Müllner, 2013), kMeans version 1.3.2 (Pedregosa et al., 2011), HDBSCAN* version 0.8.28 (McInnes et al., 2017), and FLASC version 0.1.3 with the full and core approximation graphs. HDBSCAN* and FLASC were evaluated with min samples $k=10$, min cluster size $m_{c}=100$, min branch size $m_{b}=20$, allow single cluster enabled, and their multiprocessing support disabled to better describe the algorithms’ intrinsic complexity. These parameter values allow the algorithms to find clusters and branches that are slightly smaller than how they are generated. Fastcluster’s default parameter values were used, resulting in a single linkage dendrogram. kMeans was evaluated with $k=c$, effectively attempting to recover one cluster for each starting point.
@@
Time measurements were performed with a $5.4$GHz AMD R7 7700X processor. Each algorithm was evaluated on each data set once, recording the run time and number of detected clusters. The smallest data set for which an algorithm required more than $80$ seconds was recorded for each number of dimensions. Larger data sets were not evaluated for those algorithms.
@@
Figure 6: Benchmark run times (s) over the data set size and number of dimensions. The algorithms’ scaling behaviours are visualised by quadratic regression lines relating compute time to the number of points in log–log space, shown in linear space. The shaded areas around each line indicate the regression’s $95\%$ confidence interval.
Results
@@
Figure 6 shows run times in seconds over the data set size and number of dimensions. There are three patterns of note. Firstly, fastcluster shows the steepest quadratic trend, resulting in the longest run times. FLASC and HDBSCAN* show shallower trends on the $2$ and $8$ dimensional data sets but approach fastcluster’s quadratic trend in the $16$ dimensional case. This is expected because querying space trees becomes more expensive with more dimensions (Weber et al., 1998). Secondly, HDBSCAN* and both FLASC variants scale similarly in all three conditions. HDBSCAN* is consistently the fastest of the three, followed by FLASC with the core approximation graph. However, their run time differences diminish in the higher dimensional cases, indicating that the additional cost of branch detection is relatively low compared to the cost of detecting the clusters. Finally, kMeans had the shallowest scaling trend in all three conditions, indicating that it is the fastest algorithm of the four.
@@
In conclusion, both FLASC variants’ computational performance scale similarly to HDBSCAN* and the more dimensions the data contains, the smaller the scaling trend differences between the algorithms.
Use Case: Diabetes Types
Next, we present a data exploration case in which identifying branch-based subgroups is essential to understand the data. Reaven and Miller (1979) attempted to clarify a “horse shoe”-shaped relation between glucose levels and insulin responses in diabetes patients. Three of the metabolic variables they measured turned out to be very informative in a 3D scatterplot, showing a dense core with two less-dense branches, which they considered unlikely to be a single population. Seeing that plot was instrumental in their understanding of the data (Miller, 1985).
More recently, Singh et al. (2007) used the same data set to demonstrate how Mapper with a density-based lens function visualises these flares without manually specifying which dimensions to plot. Their analysis leverages the flares’ lower density, allowing them to be detected without a centrality metric. In general, though, local density minima do not always relate to branches, especially for data sets with multiple branching clusters.
In this use case, we show how FLASC can detect the branching pattern in this data set and classify the observations by their branch without manually extracting the flares from a visualisation.
Evaluation and settings
The data set—obtained from Andrews and Herzberg (1985)—contains five variables describing 145 subjects: the relative weight, the plasma glucose level after a period of fasting, the steady-state plasma glucose response (SSPG), and two areas under a curve—one for glucose (AUGC) and one for insulin (AUIC)—representing the total amount observed during the experimental procedure described in Reaven and Miller (1979). All five variables were z-score normalised and used to compute the Euclidean distance between subjects.
Results
(a) FLASC
(b) HDBSCAN*
Figure 7: Subgroups detected by (a) FLASC and (b) HDBSCAN* shown in a 3D scatterplot over the area under the plasma glucose curve (AUGC), the area under the insulin curve (AUIC), and the steady state plasma glucose response (SSPG) from Reaven and Miller (1979). Grey points were classified as noise.
Figure 7 shows the detected subgroups encoded using colour on the 3D scatterplot. FLASC’s classification (Fig. 7(a)) distinguishes the branches from the central core. The algorithm also finds a low-persistent flare representing the central core’s bottom. This flare could be ignored by specifying a persistence threshold. In contrast, HDBSCAN*’s classification (Fig. 7(b)) does not find the branches. Instead, it finds part of the left branch as a small low-persistent cluster and merges most of the right branch with the central core.
Use Case: Cell Development
Finally, we demonstrate a use case where detecting a branch hierarchy is important for understanding the data set. Specifically, we analyse Packer et al. (2019)’s cell development atlas for the C. Elegans, a small roundworm that is often used in biological studies. They analysed gene expressions in C. Elegans embryos to uncover the trajectories along which cells develop. Broadly speaking, this data set describes what happens in cells as they develop from a single egg cell into all the different tissues that exist within fully grown C. Elegans worms.
After pre-processing, the data set appears to contain both cluster and branching structures when viewed in Packer et al. (2019)’s 3D projection. In this use case, we demonstrate that FLASC’s branch hierarchy provides interesting information about the data set’s structure even when the main subgroups can be detected as clusters.
Evaluation and settings
The data and pre-processing scripts were obtained from Monocle 3’s (Cao et al., 2019) documentation (Pliner et al., 2022). The pre-processing stages normalise the data, extract the 50 strongest PCA components, and correct for batch effects using algorithms from Haghverdi et al. (2018). HDBSCAN* was evaluated on the pre-processed data with an angular distance metric because the cosine distance metric is not supported in its optimised code path. HDBSCAN* was tuned to find multiple clusters with min samples $k=5$ and min cluster size $m_{c}=50$. FLASC was evaluated on a 3D UMAP (McInnes et al., 2020) projection that denoised the data set. Using three instead of two dimensions reduces the chance of branch overlaps in the embedding. UMAP used the angular distance metric to find $30$ nearest neighbours. The disconnection distance parameter was set to exclude the 16-percentile least dense points detected by HDBSCAN*, thereby preventing shortcuts across the data’s structure. The largest connected component in the resulting UMAP graph was projected to 3D while varying the repulsion strength to avoid crossings and ensure connected structures remained close. The same layout procedure was used to project the graph to 2D to visualise the data. FLASC was tuned to detect the branching hierarchy of the dataset as a single cluster by selecting min samples $k=3$, min cluster size $m_{c}=500$, min branch size $m_{b}=50$, enabling allow single cluster. Branches were detected using the core cluster approximation graph and selected with the leaf strategy.
Results
(a) HDBSCAN*
(b) FLASC
(c) Cluster hierarchy
(d) Branch hierarchy
Figure 8: Results for the single cell gene expression use case using 50 dimensional pre-processed data from Packer et al. (2019). (a) 2D UMAP projection (McInnes et al., 2020) coloured by HDBSCAN* clusters detected in the pre-processed data. (b) The same projection coloured by FLASC branches detected from a 3D UMAP projection. (c) and (d) show cluster and branch hierarchies, respectively. The icicle plots were adapted from McInnes et al. (2017) to indicate selected clusters with colour and labels.
Figures 8(b) and 8(a) show the 2D projections. Data points are coloured to indicate the detected clusters and branches, respectively. There are two main differences between the branches and the clusters. Firstly, two regions where branches merge are detected as clusters, namely clusters 10 and 14. These regions do not have a distinct branch label, but are identifiable in the branch hierarchy (Fig. 8(d)). Secondly, all but one branch—branch 13—are also detected as a cluster in this dataset, indicating the branches represent regions with high local density. Considering that the branches correspond to developmental end-states, it is unsurprising that local density maxima occur within them. One could imagine that the variation in gene expression is higher during development and that fully developed cells are observed more frequently. Both scenarios could cause these local density maxima.
More interesting are the differences between the cluster and branch hierarchies. Figures 8(c) and 8(d) visualise these hierarchies as an icicle plot. These designs are adapted from McInnes et al. (2017) to indicate the selected branches and clusters using colour and a label. Segment widths encode the number of points in the tree below the segment. The hierarchies highlight that while HDBSCAN* detects the branches as clusters, it does not capture the trajectories. For example, the hierarchy does not reflect that clusters 0 and 1 connect to the whole data set through cluster 10. FLASC’s branch hierarchy, on the other hand, more closely resembles the data’s shape. For example, the hierarchy describes the embedding’s left and right side with 5 and 6 branches, respectively. In general, branches that merge into the cluster near each other are also close in the branch condensed tree. For branches connected to multiple other branches in the cluster approximation graph, only the most eccentric connection is captured by the branch-condensed tree.
All in all, this use case demonstrated that FLASC’s branch hierarchy provides information about the data’s shape that may not be obvious from cluster hierarchies. In addition, we found it beneficial to suppress noisy connectivity using dimensionality reduction techniques when detecting branches.
Discussion
Two synthetic benchmarks and two real-world use cases were performed to demonstrate FLASC and its properties. We start our discussion by providing remarks for each benchmark and use case in order.
The stability benchmark compared FLASC and kMeans in their ability to detect branches that do not contain a density maximum. The benchmark quantified performance in terms of ARI and a centroid spread measure that describes output similarity between multiple samples of the same underlying distribution. While both algorithms were able to detect the branches, FLASC performed better than kMeans on both measures. In addition, kMeans requires specifying the correct number of subgroups a priori, which is difficult with unfamiliar data. This analysis could be expanded to investigate how well FLASC deals with unequal branch lengths. The weighted average data point—and eccentricity measure as a result—may not accurately reflect the centre of such clusters. Consequently, FLASC’s branch hierarchy will be a less accurate representation of the underlying topology but should still detect the branches. Monocle 3 (Qiu et al., 2017) deals with this problem by selecting the centre point in a projection manually (Pliner et al., 2022). Other eccentricity measures discussed below could also improve FLASC’s performance in such cases.
The computational performance benchmark demonstrated that FLASC’s computational performance scales similar to HDBSCAN*. The scaling trends appear to become more similar as the data contained more dimensions. Neither algorithm was evaluated with multiprocessing enabled, which can introduce run-time differences in practical applications. In addition, extracting the full cluster approximation graph can be more expensive than reported, depending on the data’s characteristics. kMeans provided even quicker run times, but is limited in usability due to lower stability and its predefined number of clusters.
The diabetes types use case demonstrated a real-world dataset in which branches that are not detectable as density-based clusters represent meaningful subgroups. FLASC is designed to detect such branches without knowing they exist a priori or extracting them from a visualisation manually. The cell development use case showed how FLASC behaves on a more complex data set. Here, the subgroups were detectable by both FLASC and HDBSCAN*. FLASC still provides a benefit for exploration because its branch hierarchy more closely resembles the data’s shape. Structure learning algorithms, as described in Section “Related Work – Structure Learning”, can provide even more information about the data’s shape at more computational costs.
FLASC’s practical value
@@ Discussion
As demonstrated by the cell development use case, the argument that branches are not detectable as clusters only applies when they do not contain local density maxima. Subpopulations tend to have some location in feature space where observations are more likely. These locations are detectable as local density maxima, allowing data points surrounding them to be classified as a particular cluster. If one is only interested in the existence of subgroups, then FLASC only provides a benefit on datasets where relevant subgroups are sparse (e.g., the diabetes types use case). If one is also interested in the clusters’ shapes, then FLASC’s branch hierarchy provides information that cannot be extracted from a cluster hierarchy. We envision FLASC as a valuable tool for exploring unfamiliar data, providing guidance into which subpopulations exist and informing follow-up questions. Knowing that a cluster may represent multiple subgroups can be very relevant.
Alternative eccentricity metrics
@@
The presented FLASC algorithm uses a geometric distance-to-centroid metric to describe how eccentric data points are within a cluster (Fig. 2(a), Eq. 3). An interesting alternative is an unweighted geodesic eccentricity, which measures path-lengths between each data point and the cluster’s root point in the cluster approximation graph. Here, the root point can be chosen as the data point closest to the cluster’s centroid, as we did for the branch-membership vectors (Fig. 3). Such a geodesic eccentricity would agree with the notion that distances in high dimensional data may not accurately reflect distances along the intrinsic structure of a data set, which was one of the motivations for Reversed Graph Embeddings (Mao et al., 2017). It would also be closer to the maximum shortest-path centrality metric used by (Vandaele et al., 2021).
Several trade-offs between the geometric and geodesic eccentricity metrics made us choose the geometric one:
  • Computing the geodesic eccentricity is more expensive because it requires an additional traversal over the entire cluster approximation graph. The extra cost, however, should be low compared to other parts of the algorithm.
  • The resolution of the unweighted geodesic eccentricity is lower, as it expresses the number of edges to the root point. As a result, zero-persistent branches are more likely to occur. In addition, it reduces the detectability of small branches that are well connected. On the other hand, that can be seen as beneficial noise suppression. In addition, the branch selection persistence parameter becomes more interpretable and would represent the traversal depth of a branch in the approximation graph.
@@
  • The cluster’s centroid may lie outside of the cluster itself, resulting in a root point and eccentricity values that do not accurately describe its centre. For example, imagine a U-shaped cluster. The centroid would lie in between the two arms, and the root would lie in one of the arms. As a result, the geodesic metric would find one smaller and one larger branch rather than two equal branches. On the other hand, the geometric eccentricity finds the two branches and the connecting bend as three separate groups. Confusingly, it also contains two regions with a local eccentricity maximum, which FLASC gives a single label. Placing the root at an arbitrary eccentric location as in Ge et al. (2011) avoids this issue but necessitates a different interpretation of the branch hierarchy and branch probability.
@@
FLASC’s general process can also be used with metrics that capture other aspects than eccentricity. At its core, FLASC consists of two filtrations, one to determine the connectivity between data points and one to analyse a signal on the resulting graph. The process would then describe how many distinct local minima (or maxima) of the metric exist within the clusters. The resulting interpretation does not have to relate to the cluster’s shape.
@@
One could even interpret FLASC as two applications of HDBSCAN*: one over the density and one over the eccentricity. This perspective raises a possible improvement to the algorithm by translating the mutual reachability concept to the centrality metric. The idea of ‘pushing away points in low-density regions’ can also be applied to the centrality and would emphasise the centrality difference between the centre and branch ends. Additionally, smoothing the centrality profile by incorporating neighbouring values could improve the algorithm’s robustness to noise. The additional computational cost should be low, as points’ neighbours are already known when the centrality is computed. Another way to improve noise robustness could be to implement the mutual $k$-nearest neighbour approach used by Dalmia and Sia (2021) to improve UMAP projections. It would provide a subgraph of the core approximation graph that better reflects the cluster’s connectivity in high dimensional data sets. We leave evaluating these ideas for future work.
Visually summarising data’s shape
@@
A strength of Mapper (Singh et al., 2007) and Reversed Graph Embeddings (Mao et al., 2017) is that they can summarise the data’s shape using intuitive visualisations. While FLASC’s branch-condensed tree provides some information about the clusters’ shapes, interpreting the shape is not trivial. Studying how well two-dimensional layouts of FLASC’s cluster approximation graphs work as shape summarising visualisations would be an interesting future research direction. These graphs directly encode the connectivity used by the algorithm. Another benefit is that—unlike in Mapper—all (non-noise) data points are represented in the graphs once. Directly visualising the graphs, however, probably does not scale to larger sizes in terms of computational cost for the layout algorithm and visual interpretability. Ways to summarise the networks would have to be found, which could be based on kMeans centroids like in Reversed Graph Embeddings (Mao et al., 2017), local density maxima in the cluster, or a Reeb-Graph approach similar to Ge et al. (2011).
Conclusion